OpenStudy (bmk614):

If dy/dx=2xy, then d2y/dx2=

OpenStudy (idku):

\(\displaystyle\frac{dy}{dx}=2xy\)

OpenStudy (idku):

Can you do the product rule on this?

OpenStudy (bmk614):

Yes. I did that and got \[2y+2x \frac{ dy }{ dx }\]

OpenStudy (bmk614):

and then I plugged that in

OpenStudy (idku):

Differentiate both sides like you normally would ... Product side for the left side (with a chain rule of y' (or dy/dx) for every derivative of y), and the right side, when differentiate, is just a second derivative. Note that your second derivative is in terms of the first derivative, but you already know that your first derivative is y'=2xy, so plug that in as well.

OpenStudy (idku):

Product rule, not product side , sorry.

OpenStudy (bmk614):

\[2y+4x^2y\]

OpenStudy (phi):

yes

OpenStudy (idku):

\(\displaystyle\frac{dy}{dx}=2xy\) \(\displaystyle\frac{d^2y}{dx^2}=2y+2x\frac{dy}{dx}\) \(\displaystyle\frac{d^2y}{dx^2}=2y+2x(2xy)=?\)

OpenStudy (bmk614):

But the answer is\[4x^2y\]

OpenStudy (idku):

No it is not, unless the problem is different.

OpenStudy (bmk614):

No that is the problem. But I looked at the answer key and that was the answer.

OpenStudy (idku):

well, your answer key is wrong then

OpenStudy (bmk614):

So the company that made the book and the answer key are wrong?

OpenStudy (idku):

all I can say is that 4x^2y is for sure not the answer. The answer is definitely, 2y+2x^2y.

OpenStudy (bmk614):

well that is what I thought too, but apparently not.

OpenStudy (idku):

Ask your teacher...

OpenStudy (idku):

and again the answer key is for sure wrong.

OpenStudy (idku):

You can ask literally anyone who knows math online, offline, in hell or in paradise.

OpenStudy (idku):

Again, \(\displaystyle \frac{dy}{dx}=2xy\) \(\displaystyle \frac{d^2y}{dx^2}=2y+2x\frac{dy}{dx}=2x+2y(2xy)=2x+4x^2y.\) Good luck!

OpenStudy (bmk614):

Chill, I agree with you.

OpenStudy (idku):

ok